# Low-temperature scaling laws in unconventional flat-band superconductors

### v1

@@ IV.2 Superfluid weight
We assume that the single-particle Hamiltonian is TRS and $N=-\partial\Omega/\partial\mu$ is constant in $\mathbf{q}$. Then, the superfluid weight is defined as the second total derivative of the free energy with respect to the external gauge field $\mathbf{q}=\mathbf{A}$ [38, 39]. For unconventional pairings, it has been shown that the superfluid weight is given by [22, 23]
@@
The first term is the conventional contribution and depends only on the curvature of the energy bands and the second term is the local part of the geometrical contribution. Analytical expressions for these contributions can be found, for example, in the supplementary material of Ref. [31]. The third term represents the functional contribution (or the nonlocal part of the geometrical contribution) to the superfluid weight. For zero temperature, an analytical expression was derived in Ref. [23]. In Appendix A, we provide a generalization of the formula for nonzero temperature (see also Refs. [22, 20] for related elaborations). Note that for conventional pairings, we can find a basis in which the functional contribution vanishes [21]. Since we consider flat-band superconductors, the conventional contribution vanishes, so that we only need to find the low-temperature scaling behaviors of the geometrical and functional contributions. Appendix B collects explicit formulas for the superfluid weight in the isolated-band limit relevant to our discussion.
@@ IV.2.1 Geometrical superfluid weight
$D_{\mathrm{s},ij}^{\mathrm{geom}}(T)\approx\frac{1}{V}\sum_{\mathbf{k}}\frac{n_{\mathrm{F}}(E_{\mathbf{k}})}{E_{\mathbf{k}}}|f(\mathbf{k})|^{2}g_{ij}(\mathbf{k})\,$, (44)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial T}=\frac{1}{T^{2}}\int\!\frac{\mathrm{d}^{2}k}{(2\pi)^{2}}g_{ij}(\mathbf{k})\frac{e^{E_{\mathbf{k}}/T}E_{\mathbf{k}}^{2}}{(e^{E_{\mathbf{k}}/T}+1)^{2}}\,$, (45)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial T}=\int_{0}^{\infty}\mathrm{d}E\,D(E)\left(\frac{E}{T}\right)^{2}\frac{e^{E/T}}{(e^{E/T}+1)^{2}}\langle g_{ij}\rangle_{E}\,$, (46)
@@
$\frac{\partial(\Delta D^{\mathrm{geom}}_{\mathrm{s},ij})}{\partial T}=\langle g_{ij}\rangle_{E=0}\int_{0}^{\infty}\mathrm{d}x\frac{x^{2}e^{x}}{(e^{x}+1)^{2}}D(xT)T\,.$ (50)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial\Delta_{T}}=\frac{\langle g_{ij}\rangle_{E=0}}{\Delta_{0}}\int_{0}^{\infty}\mathrm{d}x\,\frac{x[e^{x}(1-x)+1]}{(e^{x}+1)^{2}}D(xT)T^{2}\,.$ (51)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial T}$ $=D_{0}\langle g_{ij}\rangle_{E=0}\Gamma(\alpha+2)\eta(\alpha+1)T^{\alpha}\,.$ (52)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial\Delta_{T}}$ $=-\frac{D_{0}\langle g_{ij}\rangle_{E=0}\alpha\Gamma(\alpha+1)\eta(\alpha+1)}{\Delta_{0}}T^{\alpha+1}\,.$ (53)
@@
We then suppose that the low-energy scaling law of the DOS is given by $D(E)=D_{0}E^{\alpha-1}\ln(1/E)$. This case for $\alpha=2/m$ corresponds to a crossing of two line nodes. Because of $\ln(1/(xT))=\ln(1/T)+\ln(1/x)$, we can divide the integrals in Eq. (50) and Eq. (51) into two terms each. The evaluation of the integrals containing $\ln(1/T)$ is identical to the previous case. Moreover, the second terms containing $\ln(1/x)$ are subleading because the integrals,
@@ IV.2.2 Functional superfluid weight
$D_{\mathrm{s},ij}^{\mathrm{func}}(T)\approx\frac{1}{V}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\sum_{\alpha,\beta=1}^{N_{\mathrm{B}}}\sum_{\mu,\nu\in\left\{\mathrm{R},\mathrm{I}\right\}}\frac{M_{\alpha\mu,\beta\nu}^{-1}(\mathbf{k},\mathbf{k}^{\prime})f^{\mu}(\mathbf{k})f^{\nu}(\mathbf{k}^{\prime})n_{\mathrm{F}}(E_{\mathbf{k}})n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}})}{E_{\mathbf{k}}E_{\mathbf{k}^{\prime}}}h_{ij,\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime})\,$, (59)
where $f^{\mathrm{R}}=\mathrm{Re}(f)$ and $f^{\mathrm{I}}=\mathrm{Im}(f)$ indicate the real and imaginary parts, $h_{ij,\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime})$ is a nonlocal multi-state quantum geometric quantity defined in Eq. (113), and the components of the matrix $M^{-1}$ are given by
$M^{-1}_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})$ $=\frac{U_{0}}{V\Delta_{T}^{2}}\mathrm{Re}\big(f(\mathbf{k})f^{\ast}(\mathbf{k}^{\prime})\big)\left(\delta_{\alpha\beta}\delta_{\mu\nu}+\mathcal{O}(T^{1+\gamma})\right)$ (60)
with some exponent $\gamma>0$ depending on the node type and determined by the temperature behavior of Eq. (98). The leading order is thus provided by
$D_{\mathrm{s},ij}^{\mathrm{func}}(T)$ $=\frac{U_{0}}{V^{2}\Delta_{0}^{2}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}E_{\mathbf{k}}E_{\mathbf{k}^{\prime}}n_{\mathrm{F}}(E_{\mathbf{k}})n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}})\cos^{2}(\phi(\mathbf{k})-\phi(\mathbf{k}^{\prime}))h_{ij}(\mathbf{k},\mathbf{k}^{\prime})\,$, (61)
where $\phi(\mathbf{k})=\mathrm{arg}(f(\mathbf{k}))$ and $h_{ij}\equiv\sum_{\alpha}h_{ij,\alpha\alpha}$. Analogously to the geometrical superfluid weight, we define
$\Delta D_{\mathrm{s},ij}^{\mathrm{func}}=D_{\mathrm{s},ij}^{\mathrm{func}}(0)-D_{\mathrm{s},ij}^{\mathrm{func}}(T)$ (62)
and take the derivative with respect to the temperature to obtain
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=\frac{U_{0}}{V^{2}\Delta_{0}^{2}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\frac{E_{\mathbf{k}}E_{\mathbf{k}^{\prime}}}{T^{2}}\frac{E_{\mathbf{k}}(e^{E_{\mathbf{k}}/T}+1)e^{E_{\mathbf{k}^{\prime}}/T}+E_{\mathbf{k}^{\prime}}(e^{E_{\mathbf{k}^{\prime}}/T}+1)e^{E_{\mathbf{k}}/T}}{(e^{E_{\mathbf{k}}/T}+1)^{2}(e^{E_{\mathbf{k}^{\prime}}/T}+1)^{2}}\cos^{2}(\phi(\mathbf{k})-\phi(\mathbf{k}^{\prime}))h_{ij}(\mathbf{k},\mathbf{k}^{\prime})\,.$ (64)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=\frac{U_{0}}{\Delta_{0}^{2}}\int_{0}^{\infty}\mathrm{d}E\,\int_{0}^{\infty}\mathrm{d}E^{\prime}\frac{D(E)D(E^{\prime})EE^{\prime}}{T^{2}}\frac{E(e^{E/T}+1)e^{E^{\prime}/T}+E^{\prime}(e^{E^{\prime}/T}+1)e^{E/T}}{(e^{E/T}+1)^{2}(e^{E^{\prime}/T}+1)^{2}}\langle h_{ij}\rangle_{E,E^{\prime}}\,$, (65)
@@
$\langle h_{ij}\rangle_{E,E^{\prime}}$ $=\frac{1}{D(E)D(E^{\prime})}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\delta(E-E_{\mathbf{k}})\delta(E^{\prime}-E_{\mathbf{k}^{\prime}})\cos^{2}(\phi(\mathbf{k})-\phi(\mathbf{k}^{\prime}))h_{ij}(\mathbf{k},\mathbf{k}^{\prime})$ (66)
$=\langle h_{ij}\rangle_{E=0,E^{\prime}=0}+\mathcal{O}(|\mathbf{k}|,|\mathbf{k}^{\prime}|)\,.$ (67)
We substitute $x=E/T$ and $y=E^{\prime}/T$,
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=\frac{U_{0}\langle h_{ij}\rangle_{E=0,E^{\prime}=0}}{\Delta_{0}^{2}}\int_{0}^{\infty}\mathrm{d}x\,\int_{0}^{\infty}\mathrm{d}y\,\frac{xy[x(e^{x}+1)e^{y}+y(e^{y}+1)e^{x}]}{(e^{x}+1)^{2}(e^{y}+1)^{2}}D(xT)D(yT)T^{3}\,$, (68)
and then insert $D(E)=D_{0}E^{\alpha-1}\ln^{\beta}(1/E)$. Accordingly, the leading order is given by
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=\frac{2U_{0}K_{1,\alpha}\langle h_{ij}\rangle_{E=0,E^{\prime}=0}}{\Delta_{0}^{2}}T^{2\alpha+1}\ln^{2\beta}(1/T)\,$, (69)
where
$K_{1,\alpha}$ $=\Gamma(\alpha+2)\Gamma(\alpha+1)\eta(\alpha+2)\eta(\alpha)\,.$ (70)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial\Delta_{T}}=\frac{2U_{0}\langle h_{ij}\rangle_{E,E^{\prime}=0}}{\Delta_{0}^{3}}$
$\times\int_{0}^{\infty}\mathrm{d}x\int_{0}^{\infty}\mathrm{d}y\,\frac{xy[1+e^{x}(1-x)]}{(e^{x}+1)^{2}(e^{y}+1)}D(xT)D(yT)T^{4}.$ (71)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial\Delta_{T}}\frac{\mathrm{d}\Delta_{T}}{\mathrm{d}T}=\frac{2U_{0}K_{2,\alpha}\langle h_{ij}\rangle_{E,E^{\prime}=0}}{\Delta_{0}^{3}}T^{3\alpha+2}\ln^{3\beta}(1/T)\,$, (72)
where
$K_{2,\alpha}=-\alpha[\Gamma(\alpha+1)\eta(\alpha+1)]^{2}\,.$ (73)
@@
$\Delta D_{\mathrm{s},ij}^{\mathrm{func}}\propto T^{2\alpha+2}\ln^{2\beta}(1/T)\quad(T\ll T_{\mathrm{c}})\,.$ (74)

### v2

@@ IV.2 Superfluid weight
The superfluid weight is defined by the second total derivative of the free energy with respect to the external gauge field $\mathbf{q}=\mathbf{A}$ as [21, 49]
${D_{\mathrm{s},ij}(T)=\frac{1}{V}\frac{\mathrm{d}^{2}F}{\mathrm{d}q_{i}\mathrm{d}q_{j}}\biggl|_{\mathbf{q}=0}}\,.$ (41)
We assume that the single-particle Hamiltonian is TRS and $N=-\partial\Omega/\partial\mu$ is constant in $\mathbf{q}$. Then, for unconventional pairings, it has been shown that the superfluid weight is given by [29, 9]
@@
The first term is the conventional contribution and depends only on the curvature of the energy bands and the second term is the local part of the geometrical contribution. Analytical expressions for these contributions can be found, for example, in the Supplemental Material of Ref. [59]. The third term represents the functional contribution (or the nonlocal part of the geometrical contribution) to the superfluid weight. For zero temperature, an analytical expression was derived in Ref. [9]. Note that for conventional pairings, we can find a basis in which the functional contribution vanishes [22]. Since we consider flat-band superconductors, the conventional contribution vanishes, so that we only need to find the low-temperature scaling behaviors of the geometrical and functional contributions. Appendix A collects explicit formulas for the superfluid weight in the isolated-band limit relevant to our discussion.
@@ IV.2.1 Geometrical superfluid weight
$\Delta D_{\mathrm{s},ij}^{\mathrm{geom}}\approx\frac{8}{V}\sum_{\mathbf{k}}\frac{n_{\mathrm{F}}(E_{\mathbf{k}})}{E_{\mathbf{k}}}|f(\mathbf{k})|^{2}g_{ij}(\mathbf{k})\,$, (45)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial T}=\frac{8}{T^{2}}\int\!\frac{\mathrm{d}^{2}k}{(2\pi)^{2}}g_{ij}(\mathbf{k})\frac{e^{E_{\mathbf{k}}/T}E_{\mathbf{k}}^{2}}{(e^{E_{\mathbf{k}}/T}+1)^{2}}\,$, (46)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial T}=8\int_{0}^{\infty}\mathrm{d}E\,D(E)\left(\frac{E}{T}\right)^{2}\frac{e^{E/T}}{(e^{E/T}+1)^{2}}\langle g_{ij}\rangle_{E}\,$, (47)
@@
$\frac{\partial(\Delta D^{\mathrm{geom}}_{\mathrm{s},ij})}{\partial T}=8\langle g_{ij}\rangle_{E=0}\int_{0}^{\infty}\mathrm{d}x\frac{x^{2}e^{x}}{(e^{x}+1)^{2}}D(xT)T\,.$ (51)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial\Delta_{T}}=\frac{8\langle g_{ij}\rangle_{E=0}}{\Delta_{0}}\int_{0}^{\infty}\mathrm{d}x\,\frac{x[e^{x}(1-x)+1]}{(e^{x}+1)^{2}}D(xT)T^{2}\,.$ (52)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial T}$ $=8D_{0}\langle g_{ij}\rangle_{E=0}\Gamma(\alpha+2)\eta(\alpha+1)T^{\alpha}\,.$ (53)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{geom}})}{\partial\Delta_{T}}$ $=-\frac{8D_{0}\langle g_{ij}\rangle_{E=0}\alpha\Gamma(\alpha+1)\eta(\alpha+1)}{\Delta_{0}}T^{\alpha+1}\,.$ (54)
@@
We then suppose that the low-energy scaling law of the DOS is given by $D(E)=D_{0}E^{\alpha-1}\ln(1/E)$. This case for $\alpha=2/m$ corresponds to a crossing of two line nodes. Because of $\ln[1/(xT)]=\ln(1/T)+\ln(1/x)$, we can divide the integrals in Eq. (51) and Eq. (52) into two terms each. The evaluation of the integrals containing $\ln(1/T)$ is identical to the previous case. Moreover, the second terms containing $\ln(1/x)$ are subleading because the integrals,
@@ IV.2.2 Functional superfluid weight
Analogously to the geometrical superfluid weight, we define
$\Delta D_{\mathrm{s},ij}^{\mathrm{func}}=D_{\mathrm{s},ij}^{\mathrm{func}}(0)-D_{\mathrm{s},ij}^{\mathrm{func}}(T)\,.$ (60)
@@
$\Delta D_{\mathrm{s},ij}^{\mathrm{func}}\approx$ $\frac{4}{V}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\sum_{\alpha,\beta=1}^{N_{\mathrm{B}}}\sum_{\mu,\nu\in\left\{\mathrm{R},\mathrm{I}\right\}}M_{\alpha\mu,\beta\nu}^{-1}(\mathbf{k},\mathbf{k}^{\prime})f^{\bar{\mu}}(\mathbf{k})f^{\bar{\nu}}(\mathbf{k}^{\prime})$
$\times{s_{\mu}s_{\nu}}\frac{{2(n_{\mathrm{F}}(E_{\mathbf{k}})+n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}}))}{-4}n_{\mathrm{F}}(E_{\mathbf{k}})n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}})}{E_{\mathbf{k}}E_{\mathbf{k}^{\prime}}}$
$\times h_{ij,\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime}),$ (61)
where we define $s_{\mathrm{R},\mathrm{I}}=\pm 1$, $f^{\mathrm{R}}=\mathrm{Re}(f)$ and $f^{\mathrm{I}}=\mathrm{Im}(f)$ indicate the real and imaginary parts, $\bar{\mu}=\mathrm{R}$ if $\mu=\mathrm{I}$ and $\bar{\mu}=\mathrm{I}$ if $\mu=\mathrm{R}$, $h_{ij,\alpha\beta}(\mathbf{k},\mathbf{k}^{\prime})$ is a nonlocal multi state quantum geometric quantity defined in Eq. (93), and the components of the matrix $M^{-1}$ are given by
$M^{-1}_{\alpha\mu,\beta\nu}(\mathbf{k},\mathbf{k}^{\prime})=$ $\frac{1}{V}\!\left\{{\mathrm{Re}[U(\mathbf{k},\mathbf{k}^{\prime})]\delta_{\mu\nu}-\mathrm{Im}[U(\mathbf{k},\mathbf{k}^{\prime})]\epsilon_{\mu\nu}}\right\}\!\delta_{\alpha\beta}$
$+\mathcal{O}(T^{\gamma})$ (62)
with some exponent $\gamma>0$ depending on the node type and determined by Eq. (97). Hence, at leading order in temperature, the functional contribution to the superfluid weight is given by
$\Delta D_{\mathrm{s},ij}^{\mathrm{func}}\approx$ $\frac{4}{V^{2}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\left[2(n_{\mathrm{F}}(E_{\mathbf{k}})+n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}}))-4n_{\mathrm{F}}(E_{\mathbf{k}})n_{\mathrm{F}}(E_{\mathbf{k}^{\prime}})\right]$
$\times\mathrm{Re}\!\left[U(\mathbf{k},\mathbf{k}^{\prime})e^{i[\phi(\mathbf{k}^{\prime})-\phi(\mathbf{k})]}\right]h_{ij}(\mathbf{k},\mathbf{k}^{\prime})\,$, (63)
where $\phi(\mathbf{k})=\mathrm{arg}(f(\mathbf{k}))$ and $h_{ij}\equiv\sum_{\alpha}h_{ij,\alpha\alpha}$. We take the derivative with respect to the temperature to obtain
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=\frac{8}{V^{2}}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\frac{1}{T^{2}}\frac{E_{\mathbf{k}}e^{E_{\mathbf{k}}/T}(e^{2E_{\mathbf{k}^{\prime}}/T}-1)+E_{\mathbf{k}^{\prime}}e^{E_{\mathbf{k}^{\prime}}/T}(e^{2E_{\mathbf{k}}/T}-1)}{(e^{E_{\mathbf{k}}/T}+1)^{2}(e^{E_{\mathbf{k}^{\prime}}/T}+1)^{2}}{\mathrm{Re}\!\left[U(\mathbf{k},\mathbf{k}^{\prime})e^{i[\phi(\mathbf{k}^{\prime})-\phi(\mathbf{k})]}\right]}h_{ij}(\mathbf{k},\mathbf{k}^{\prime})\,.$ (65)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=8\int_{E,E^{\prime}=0}^{\infty}\frac{D(E)D(E^{\prime})}{T^{2}}\frac{Ee^{E/T}(e^{2E^{\prime}/T}-1)+E^{\prime}e^{E^{\prime}/T}(e^{2E/T}-1)}{(e^{E/T}+1)^{2}(e^{E^{\prime}/T}+1)^{2}}\langle{\mathrm{Re}\!\left[U(\mathbf{k},\mathbf{k}^{\prime})e^{i[\phi(\mathbf{k}^{\prime})-\phi(\mathbf{k})]}\right]}h_{ij}(\mathbf{k},\mathbf{k}^{\prime})\rangle_{E,E^{\prime}},$ (66)
@@
$\langle F\rangle_{E,E^{\prime}}$ $=\frac{1}{D(E)D(E^{\prime})}\sum_{\mathbf{k},\mathbf{k}^{\prime}}\delta(E-E_{\mathbf{k}})\delta(E^{\prime}-E_{\mathbf{k}^{\prime}})F(\mathbf{k},\mathbf{k}^{\prime})\,.$ (67)
Since the basis functions are analytic functions, the pairing potential admits the Taylor expansion,
$U(\mathbf{k},\mathbf{k}^{\prime})=U_{0}\sum_{i=1}^{d_{\Gamma}}\sum_{M,N\geq 0}\frac{1}{M!N!}\left(\partial^{M}\bar{\psi}_{i}^{\Gamma}(0)\right)\left(\partial^{N}\psi_{i}^{\Gamma}(0)\right)\mathbf{k}^{M}(\mathbf{k}^{\prime})^{N}\,$, (68)
where $M,N\in\mathbb{N}^{2}_{0}$ are multi-indices. We then denote the lowest non-vanishing order of $U(\mathbf{k},\mathbf{k}^{\prime})$ by $M_{0},N_{0}$ and define the superconducting state dependent exponent $b\geq 0$ via the equation,
$\frac{U_{0}}{M_{0}!N_{0}!}\sum_{i=1}^{d_{\Gamma}}\mathrm{Re}\!\left[\left(\partial^{M_{0}}\bar{\psi}_{i}^{\Gamma}(0)\right)\left(\partial^{N_{0}}\psi_{i}^{\Gamma}(0)\right)\Big\langle e^{i[\phi(\mathbf{k}^{\prime})-\phi(\mathbf{k})]}\mathbf{k}^{M_{0}}(\mathbf{k}^{\prime})^{N_{0}}\Big\rangle_{E,E^{\prime}}\right]=c(EE^{\prime})^{b}\,$, (69)
where the proportionality coefficient $c$ depends on the order parameter at zero temperature. According to Eq. (7), we always have $b=1$ for one-dimensional superconducting states with $d_{\Gamma}=1$. Moreover, the same holds for all two-dimensional cases considered below in Sec. V. We insert this expansion and substitute $x=E/T$ and $y=E^{\prime}/T$ in Eq. (66),
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=8c\langle h_{ij}\rangle_{E,E^{\prime}=0}\int_{x,y=0}^{\infty}\frac{x^{{b}}y^{{b}}[xe^{x}(e^{2y}-1)+ye^{y}(e^{2x}-1)]}{(e^{x}+1)^{2}(e^{y}+1)^{2}}D(xT)D(yT)T^{{1+2b}}\,$, (70)
and then insert $D(E)=D_{0}E^{\alpha-1}\ln^{\beta}(1/E)$. Since these integrals are divergent, we introduce an energy cutoff $\Lambda>0$ and find the leading order to be given by
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial T}=16D_{0}^{2}c\langle h_{ij}\rangle_{E,E^{\prime}=0}K_{\alpha\beta b}T^{{\alpha+b-1}}\ln^{\beta}(1/T),$ (71)
where the Feynman trick and the fact that the incomplete Gamma functions obey the property $\Gamma(n,x)+\gamma(n,x)=\Gamma(n)$, cf. [16, §8.35], is used to obtain
$K_{\alpha\beta b}=(-1)^{\beta}\eta(\alpha+b)\Gamma(\alpha+b+1)\frac{\mathrm{d}^{\beta}}{\mathrm{d}s^{\beta}}\biggl|_{s=0}\frac{\Lambda^{\alpha+b-s}}{\alpha+b-s}\,.$ (72)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial\Delta_{T}}=$ $-\frac{16c\langle h_{ij}\rangle_{E,E^{\prime}=0}}{\Delta_{0}}\int_{x,y=0}^{\infty}\frac{x^{{b}}y^{{b}}[xe^{x}(e^{y}-1)]}{(e^{x}+1)^{2}(e^{y}+1)}$
$\times D(xT)D(yT)T^{{2+2b}}.$ (73)
@@
$\frac{\partial(\Delta D_{\mathrm{s},ij}^{\mathrm{func}})}{\partial\Delta_{T}}\frac{\mathrm{d}\Delta_{T}}{\mathrm{d}T}=$ $\frac{16D_{0}^{2}c\langle h_{ij}\rangle_{E,E^{\prime}=0}K_{\alpha\beta b}}{\Delta_{0}}$
$\times T^{{2\alpha+b}}\ln^{{2\beta}}(1/T)\,$, (74)
@@
$\Delta D_{\mathrm{s},ij}^{\mathrm{func}}\propto T^{{\alpha+b}}\ln^{{\beta}}(1/T)\quad(T\ll T_{\mathrm{c}})\,.$ (75)